1.    If water is cooled from 50OC to -5OC, what is the temperature change on the Celsius, Fahrenheit and Rankine scales?

1
Expert's answer
2021-02-17T10:39:52-0500

Change in celsius, "\Delta T = 50-(-5) = 55^\circ C"


Initial temperature in Fahrenheit, "T_1 = \frac{(50× 9)}{5} + 32 = 122 F"

final temperature in Fahrenheit, "T_2 = \frac{(-5× 9)}{5} + 32 = 23 F"

Change in Fahrenhiet, "\Delta T = 122-23 = 99^\circ F"


Initial Temperature in Rankine, "T_1 = \frac{50 × 9}{5} + 491.67 = 581.67 R"

final Temperature in Rankine, "T_2 = \frac{-5 × 9}{5} + 491.67 = 482.67 R"

Change in temperature, "\Delta T = 581.67-482.67 = 99^\circ R"